OpenStudy (anonymous):

simplify the complex rational expression 6x+ 3/y divided by 9x+3/y

OpenStudy (anonymous):

I got (2x+1)/(3x+1)

OpenStudy (anonymous):

\[(6xy+3)/y \div (9xy+3)/y\] Take reciprocal of 9xy... and it becomes: \[(6xy+3)/y \times y/(9xy+3)\]

OpenStudy (anonymous):

ys cancel and it becomes: \[(6xy+3) \div (9xy+3)\] take 3 common and you're left with \[(2xy+1) \div (3xy+1)\]

OpenStudy (anonymous):

on the top only the 3 is divided by y. and on the bottom the whole 9x+3 is divided by y

OpenStudy (anonymous):

\[6x+(3/y)/(9x+3)/y \]

OpenStudy (anonymous):

uh oh, you should've writeen it like that..

OpenStudy (anonymous):

yeah getting it to the computer is challenging still.. I wish I could just take a snap shot. lol

OpenStudy (dtan5457):

Put this in equation form so It's easier to read.

OpenStudy (dtan5457):

Or draw it.

OpenStudy (anonymous):

\[-6/(y+2)(y+3)/[2/(y+3)]+[3/(y+2)]\] better?

OpenStudy (dtan5457):

The original question was \[\frac{ 6x+3 }{ y }\div \frac{ 9x+3 }{ y }\]

OpenStudy (dtan5457):

?

OpenStudy (anonymous):

\[\frac{ -6 }{ (x+2)(x+3) }\div \frac{ 2 }{ (y+3) }- \frac{ 3 }{ (y+2) }\] these are stacked on top of each other not written out like this

OpenStudy (anonymous):

number 8.

OpenStudy (anonymous):

for number 8 i got (2x+1) _______ (3x+1)

OpenStudy (dtan5457):

Not too sure. I'm getting\[\frac{ 6xy+3 }{ 9x+3 }\]

OpenStudy (anonymous):

was your LCD y

OpenStudy (dtan5457):

y/1 for everything

OpenStudy (anonymous):

OHHHHHHH you have to take the LCD and multiply 6 by it as well..

OpenStudy (dtan5457):

Why?

OpenStudy (anonymous):

well isnt that what you do when you multiply by the LCD? you have to multiply all parts of the equation

OpenStudy (dtan5457):

that's what i did. y/1 for everything 6x times y/1=6xy 3/y times y/1=3 9x+3/y times y/1=9x+3

OpenStudy (anonymous):

okay.. that makes sense. I now have to go back through this assignment and fix all my answers because I didn't do this. :D

OpenStudy (dtan5457):

\[\frac{ 2xy+1 }{ 3x+1 }\]

OpenStudy (dtan5457):

the most simplified answer i can get

OpenStudy (dtan5457):

after i divide my original answer by 3

OpenStudy (anonymous):

ok thanks for your help :)

OpenStudy (dtan5457):

i am fairly confident about my answer, but i hope someone checks it

OpenStudy (anonymous):

Honestly its just a homework assignment so well be going through them tomorrow am in class, so if its not right i can see why and then go from there. Im sure its right. lol your probably in a higher math class then I am anyways haha

OpenStudy (dtan5457):

im in algebra 2/trig

OpenStudy (dtan5457):

we were just learning this

OpenStudy (dtan5457):

everytime i attempt to solve problems l get 9/10 steps correct then one wrong i hope it wasn't in this case

OpenStudy (dtan5457):

@Nnesha

Nnesha (nnesha):

yes which one ?? give me question

OpenStudy (dtan5457):

scroll up there is an attachment #8

Nnesha (nnesha):

question 8 n the file right ??

OpenStudy (dtan5457):

yes

Nnesha (nnesha):

okay lets see

OpenStudy (anonymous):

thanks Nnesha

Nnesha (nnesha):

yes @dtan5457 right :)

OpenStudy (dtan5457):

wooo thank you!

Nnesha (nnesha):

yes that's suppose to be 2xy

Nnesha (nnesha):

bec  oky i just did the top part remember when you multiply variables you should add their exponents so that's why Y^2